# Spin Photonic Topological Insulator Leaky Wave Antenna By Armchair Arrangement

### v1

@@ I Introduction
Photonic topological insulators(PTIs) newly advent to use in waveguiding and antenna applications, which are robust to scattering from manufacturing defects and sharp turns[1]. One of PTI types is spin PTIs, which are reciprocal and realized by spin-degenerate metamaterials[2].
@@ III Antenna structure
Figure 5: The Scattering parameters of the proposed antenna. The yellow domains are the frequencies that the armchair part of the antenna radiates exclusively, while in the green part, both armchair and zigzag radiates. The cyan domains belong to the gap of the edge modes, where the return loss rises.
The realized gain pattern at 15.8 GHz, where radiation occurs in a broadside direction, is illustrated in Fig.6. The comparison between the patterns of the zigzag arrangement (Fig.1) and the armchair arrangement indicates that the armchair structure is more suitable than the zigzag one. However, as shown in Fig.5, the leakage factor of the zigzag structure is greater than that of the armchair structure. This weaker attenuation results in the radiation not experiencing a sudden decrease in the open stopband regions.

### v2

@@ I Introduction
Photonic topological insulators(PTIs) newly advent to use in guiding-wave and antenna applications, which are robust to scattering from manufacturing defects and sharp turns[1]. One of PTI types is spin PTIs, which are reciprocal and realized by spin-degenerate metamaterials[2].
@@ III Antenna structure
Figure 5: The Scattering parameters of the proposed antenna.The yellow areas represent the frequencies at which only the armchair portion of the antenna structure radiates. In the green area, both the armchair and zigzag parts radiate simultaneously. The cyan regions indicate the gap in the edge modes, where the return loss increases.
The realized gain pattern at 15.8 GHz, where radiation occurs in a broadside direction, is illustrated in Fig.6. The comparison between the patterns of the zigzag arrangement (Fig.1) and the armchair arrangement indicates that the armchair structure is more suitable than the zigzag one. However, as shown in Fig.5, in the green area where radiation from the zigzag section of the structure occurs, the insertion loss declines rapidly. This qualitatively indicates that the leakage factor of the zigzag structure is higher than that of the armchair structure. As a result of this weaker attenuation, the radiation does not experience a sudden drop in the open stopband regions.